OpenStudy (avountae2003):

5 1/3 + (-3 9/18) and show work please @Studentc14

OpenStudy (anonymous):

Use Please Excuse My Dear Aunt Sally PEMDAS Parenthesis Exponent Multiply Divide Add Subtract

OpenStudy (avountae2003):

can u explain more please

OpenStudy (anonymous):

you do the parenthesis first then the exponents (there are none) then multiply then divide then add then subtract

OpenStudy (avountae2003):

wat do u multiply

OpenStudy (anonymous):

in this case, you don't

OpenStudy (avountae2003):

so wat do i do

OpenStudy (anonymous):

If there is more than one thing to do, you do them in the PEMDAS order.

OpenStudy (avountae2003):

im confused

OpenStudy (anonymous):

look here. http://en.wikipedia.org/wiki/Order_of_operations

OpenStudy (avountae2003):

can u explain it for me

OpenStudy (avountae2003):

is the answer 11/6

OpenStudy (anonymous):

Exact Solution:11/6 Approximate : 1.8333 im trying to retrace my steps bein that i udes PEMDAS

OpenStudy (anonymous):

*used

OpenStudy (avountae2003):

can u show me how u got that answer

OpenStudy (anonymous):

yeah, give me a second

OpenStudy (anonymous):

Since any number divided by itself equals 1, you can rewrite 1 as a fraction, e.g. 1=2 2 =3 3 =6 6 Any whole number can be rewritten as a fraction by multiplying it by 1 in the form of a fraction. 5 can be made into a fraction like so: 5∗1=5∗3 3 =15 3 51 3 =15 3 +1 3 =15+1 3 =16 3 and there you go

OpenStudy (avountae2003):

im confused

OpenStudy (anonymous):

Try this website, it's really helpful. http://www.mathsisfun.com/operation-order-pemdas.html